Using Maxwell’s equations in free space, derive the wave equation for the z-component of
the electric field vector.
1
Expert's answer
2019-09-02T15:41:20-0400

The Maxwell's equations in free space can be written in the form as follows (Gauss units are used):


"\\nabla \\cdot \\vec{E} = \\rho,\\\\\n \t\\nabla \\cdot \\vec{H} = 0, \\\\\n \t\\nabla \\times \\vec{E} = - \\frac{1}{c} \\frac{\\partial \\vec{H}}{\\partial t},\\\\\n \t\\nabla \\times \\vec{H} = \\frac{1}{c} \\frac{\\partial \\vec{E}}{\\partial t}\t+ \\frac{4 \\pi}{c} \\vec{j}."

The curl from the left side of the third equation above is equal to:


"\nabla \times (\nabla \times \vec{E}) = \nabla(\nabla \cdot \vec{E}) - \Delta \vec{E} = \nabla \rho - \Delta \vec{E}."

Operators "\nabla" and "\frac{\partial}{\partial t}" commute, hence, calculating the curl from the right side of the same equation, we obtain:


"\\nabla \\times \\left( - \\frac{1}{c} \\frac{\\partial \\vec{H}}{\\partial t} \\right) = - \\frac{1}{c} \\frac{\\partial}{\\partial t} \\left( \\frac{1}{c} \\frac{\\partial \\vec{E}}{\\partial t}\t\n \t+ \\frac{4 \\pi}{c} \\vec{j} \\right) = - \\frac{1}{c^2} \\frac{\\partial^2 \\vec{E}}{\\partial t^2} - \\frac{4 \\pi}{c^2} \\frac{\\partial \\vec{j}}{\\partial t}"

Putting both expressions together, one can deduce:


"\Delta \vec{E} - \frac{1}{c^2} \frac{\partial^2 \vec{E}}{\partial t^2} = \nabla \rho + \frac{4 \pi}{c^2} \frac{\partial \vec{j}}{\partial t},"

which projection on the z-axis results in:


"\Delta E_z - \frac{1}{c^2} \frac{\partial^2 E_z}{\partial t^2} = \frac{\partial \rho}{\partial z} + \frac{4 \pi}{c^2} \frac{\partial j_z}{\partial t}."